# Monte Carlo simulations of relativistic shock breakout from a stellar wind

### v1

@@ 7 Summary and discussion
In this study, we performed Monte Carlo simulations of photon starved relativistic RRMS, incorporating the effect of photon escape from the upstream region under the assumption of steady-state flow in a planar geometry. These simulations are applicable to RRMS propagating through shallowly decaying density profiles, such as those found in stellar winds. The calculations were carried out for shock Lorentz factors of $\Gamma_{u}=2$, $6$, and $10$, marking the first simulation to model relativitic shock breakout emission from the first principles in the photon starved regime. [^13] Lundman & Beloborodov (2021) conducted a first principles calculation of relativistic shock breakout in the “photon rich” regime, where photon production and pair production can be neglected. We explored a wide range of energy fractions carried by escaping photons, $f_{esc}$, and analyzed the resulting shock structures and the spectral properties of the escaping radiation. We also conducted a detailed comparison between our numerical results and the shock structures predicted by an analytical model (GNL18), identifying deviations that suggest the need to reconsider the predicted observables of relativistic shock breakouts in a wind environment.
@@
    While our results confirm the overall trends predicted by the analytical model, they also reveal systematic deviations. The model successfully reproduces the structure of infinite shocks for all tested Lorentz factors ($\Gamma_{u}=2$, $6$, and $10$), and captures finite shock structures well for $\Gamma_{u}=2$ across a wide range of $f_{\rm esc}$. However, for higher Lorentz factors ($\Gamma_{u}=6$ and $10$), our simulations yield systematically narrower shock widths than predicted, particularly at large $f_{\rm esc}$. Specifically, we find a scaling of $\Delta\tilde{\tau}\propto f_{\rm esc}^{-1.6}$, compared to the analytical prediction of $\Delta\tilde{\tau}\propto f_{\rm esc}^{-1}$. Furthermore, while the analytical model predicts a positive scaling with Lorentz factor ($\Delta\tilde{\tau}\propto\Gamma_{u}$), our simulations show an inverse trend, $\Delta\tilde{\tau}\propto\Gamma_{u}^{-2}$. These discrepancies arise from a shift in the dominant deceleration mechanism at high escape fractions. In this regime, $\gamma$$\gamma$ pair production between high-energy photons—upscattered by subshock-heated pairs—and the bulk population of counterstreaming photons becomes the main source of flow deceleration, overtaking the interaction between counterstreaming photons and the advected plasma originally assumed in the analytical model.
@@
    Due to efficient temperature regulation in the downstream region, the spectra of escaping photons consistently peak at $E_{p}\approx 300$-$600~{}{\rm keV}$ in the shock frame, largely independent of $f_{\rm esc}$ and $\Gamma_{u}$. The spectral shape deviates significantly from a Wien distribution, exhibiting broadening at both low and high energies. At lower energies ($h\nu\ll E_{p}$), a flat spectral component with $f_{\nu}\propto\nu^{0}$ emerges due to free-free emission, and its prominence increases with larger $f_{esc}$. At higher energies ($h\nu>E_{p}$), our simulations reveal a spectral extension, particularly for $\Gamma_{u}=6$ and $10$, originating from inverse Compton scattering by hot pairs in the post-subshock region. The strength and extent of this high-energy component increase with both $\Gamma_{u}$—reflecting higher post-subshock temperatures—and $f_{esc}$—due to reduced optical depth for $\gamma$$\gamma$ pair production. These results indicate that the spectrum undergoes rapid broadening during breakout as $f_{\rm esc}$ increases.
@@
    Based on the deviations from the analytical model, we derive an updated closure relation applicable to relativistic breakouts with $\Gamma_{bo}\gtrsim 6$. We also present revised estimates of breakout observables in wind environments using a spherical explosion model. Notably, we find that the breakout duration becomes largely insensitive to the explosion energy and ejecta mass, depending primarily on the progenitor radius and wind optical depth, following $t_{bo}\sim 170R_{*,11}\tau_{w*}~{}{\rm s}$. This duration is significantly longer—by orders of magnitude—than that predicted by the analytical model for the same explosion parameters, reflecting the substantially reduced shock width in finite shocks. For example, a highly energetic explosion with isotropic equivalent energy $E_{ex}\sim 5\times 10^{53}{\rm erg}$ from a compact progenitor ($R_{*}\sim 10^{11}{\rm cm}$) surrounded by a wind of optical depth $\tau_{*}\sim 1$ can produce a relativistic breakout with $\Gamma_{bo}\sim 6$. The resulting emission peaks at $\sim 3kT_{obs}\sim 3~{}{\rm MeV}$, with total radiated energy $E_{bo}\sim 7\times 10^{49}~{}{\rm erg}$ and a duration of approximately $170~{}{\rm s}$ In contrast, the analytical model predicts a sub-second duration for the same parameters. Such MeV-range emission may be detectable by current gamma-ray observatories such as Fermi-GBM up to distances of $\sim$100 Mpc. Future missions with enhanced sensitivity in the MeV band, including AMEGO-X, COSI, and e-ASTROGAM, could extend this detection horizon to several hundred Mpc, significantly increasing the prospects for observing relativistic shock breakouts.
Despite the valuable insights gained from our simulations, several limitations must be acknowledged. First, as noted in the main text, our simulations cover only a limited region of the parameter space in $\Gamma_{u}$ and $f_{\rm esc}$, leaving some uncertainty in our rough estimates of the breakout emission. In addition, there are physical effects that may play an important role in RRMSs but are not included in the present study. For instance, we do not account for particle acceleration that may arise in the collisionless subshock. Furthermore, as demonstrated by Levinson (2020), strong pair enrichment in RRMSs can lead to a decoupling between protons, electrons, and positrons, where the electric fields generated by small charge separations are no longer sufficient to maintain tight coupling. This decoupling may induce plasma instabilities and lead to microturbulence, as shown in Vanthieghem et al. (2022). Such plasma effects, if present, could modify the escaping spectra through inverse Compton scattering by high-energy pairs produced by the plasma effects and may also alter the deceleration process by changing the efficiency of $\gamma$-$\gamma$ pair production within the shock. On the other hand, Mahlmann et al. (2023) showed that even a small magnetic field, neglected in our simulations, may be sufficient to maintain tight coupling between plasma species, potentially suppressing such instabilities. These issues warrant further investigation in future work.
Another key limitation is our assumption of a steady-state shock structure. In this study, we assumed that, during breakout in a shallow density gradient, the shock can adiabatically adjust its structure. However, as also discussed in GBL18, the validity of this assumption remains uncertain. Moreover, the assumption of planar geometry could also lead to deviations. Indeed, recent work by Wasserman et al. (2025), based on dynamical simulations of sub-relativistic breakouts, suggests that the planar and steady-state assumptions adopted in earlier works (Ioka et al., 2019, ILN20b) may break down in sub-relativitic shocks. We note, however, that relativistic and sub-relativistic breakouts differ substantially. In relativistic shocks, the breakout proceeds gradually with increasing $f_{\rm esc}$ due to their ability to self-generate opacity through pair production, enabling the shock to propagate over several orders of magnitude in radius during the breakout phase. In contrast, in sub-relativistic RMSs, the breakout phase cannot extend well beyond the optical depth of $\Delta\tilde{\tau}\sim\beta_{u}^{-1}$. Nevertheless, validating our assumptions will ultimately require dynamical simulations of relativistic shock breakouts—an important but challenging direction for future research.
@@ Acknowledgements
This work was supported by JSPS KAKENHI Grant Number JP24H01826, JP23K25874, and JP25H00675. S.N. is supported by the JST ASPIRE Program ”RIKEN-Berkeley mathematical quantum science initiative”. AL and EN acknoweldge support by the Israel Science Foundation grant 1995/21, and a BSF grant (2022314) and ERC grant JetNS (818899). Numerical computations and data analysis were carried out on Hokusai BigWaterfall2 system at RIKEN, XC50 at Center for Computational Astrophysics, National Astronomical Observatory of Japan, and the Yukawa Institute Computer Facility.
@@ Appendix A Spectrum at low energies
Figure 7: Shock-frame, $f_{\nu}$ flux of escaping photons normalized by the total kinetic energy flux of baryons at the upstream boundary, $F_{b}=\Gamma_{u}(\Gamma_{u}-1)n_{u}m_{p}c^{3}\beta_{u}$. The top, middle, and bottom panels display the results for shock velocities $\Gamma_{u}=2$, $6$, and $10$, respectively. Within each panel, various lines represent different escape fractions, as specified in the legend. Note that the energy range of the spectra varies across different panels.
@@
Figure 8: The low cutoff energy for free-free emission, estimated from Equation (15), is shown for $\Gamma_{u}=2$ (top), $\Gamma_{u}=6$ (middle), and $\Gamma_{u}=10$ (bottom). Within each panel, various lines represent different escape fractions, as specified in the legend.
As the equation indicates, $\nu_{cut}$ varies with position, depending on the local density of pairs and temperature. In the post-subshock region, where most photons are produced, the cutoff frequency progressively decreases in the downstream direction. The spatial profile of $\nu_{cut}$ is displayed in Figure 8. As seen in the figure, $h\nu_{cut}$ reaches maximum values of approximately $0.4$, $8$, and $30~{}{\rm eV}$ in simulations with the highest escape fractions for $\Gamma_{u}=2$, $6$, and $10$, respectively, at the immediate post-subshock region. These peak values are higher for higher escape fractions, reflecting elevated post-subshock temperatures. Beyond the peak, $h\nu_{cut}$ decreases to about $0.03$, $0.2$, and $0.3~{}{\rm eV}$ at $\tau_{*}=10$, which corresponds to roughly one diffusion length downstream from the subshock. While the maximum values somewhat depend on the escape fractions particularly for high $\Gamma_{u}$ shocks, the profile beyond $\tau_{*}\gtrsim 1$ shows little variation across different escape fractions. This stability is attributed to the minimal variation in the temperature profile that follows the rapid cooling of subshock-heated pairs (Figure 1-3).
@@ C.1 Conversion of Optical Depth
As shown in the previous section, the analytical model provides the shock profile as a function of the net optical depth, $\tau$, without explicitly specifying its relation to the spatial scale. Consequently, converting $\tau$ into a physical length scale is necessary to facilitate the comparison between the simulation and the model. In this study, we achieve this by translating $\tau$ to the pair unlodaded Thomson optical depth $\tilde{\tau}$ ($=2\Gamma_{u}\sigma_{T}z$), following the methodology outlined in GNL18:
@@
$X=\Gamma(1+a\hat{T}),$ (25)
represents an approximate value for the typical photon energy of the counterstreaming photons observed in the rest frame of the pairs, normalized by $m_{e}c^{2}$ in the case of Compton scattering. At the same time, it also represents the product of the counterstreaming photon energy and the thermal photon energy advected with the flow, both evaluated in the comoving frame. [^17] When $X\gg 1$ and the typical advected photon energy satisfies $kT\gg m_{e}c^{2}$, Equation (24) provides a good approximation for the $\gamma$-$\gamma$ pair production cross-section. There is a slight modification to the cross-section given in Equation (24) compared to GNL18, which does not include the additional $+0.5$ term in the numerator. This term is added solely to improve accuracy in the deep KN regime ($X\gg 1$), although it is not crucial, as no significant differences are observed regardless of which prescription is used. The parameter $a$, which is expected to be of order unity, is treated as an additional free parameter of the analytical model, along with $\eta$, both of which depend on the energy and angular distributions of pairs and photons. The above prescription approximates that each particle in the primary beam contributes equally to the opacity for the counterstreaming photons, regardless of whether the particle is a pair or a photon.
@@ C.2 Fitting to the simulation results
The fitting is performed by scanning the free parameters $\eta$ in the range $[0.25,1.0]$ and $a$ in the range $[0.5,3]$ with an interval of $0.05$, minimizing the squared residuals between $\log(\tilde{\tau})$ from the simulation and the analytical solution. In this section, we present the fitted results, focusing on the case of infinite shocks, i.e., $f_{esc}=0$. Figures 9-11 display the comparison of the simulation results with the best-fit analytical model. The figures show not only the $\Gamma$ profiles, which were used to optimize the fitting, but also the profiles of $\hat{T}$ and $x_{l}$. Note that $x_{l}$, a quantity introduced in the analytical model, includes contributions from both advected photons and produced pairs. However, the analytical model does not distinguish between these components and therefore does not provide their relative abundances. In contrast, in the simulations, $x_{l}$ is calculated by evaluating the individual components, $n_{\pm}$ and $n_{\gamma\rightarrow d}$, and substituting them into Equation (17). Note that $n_{\gamma\rightarrow d}$ in the numerical simulations does not include only the backscattered photons which was originally counterstreaming photons before the scattering, but also includes the photons produced by the thermal bremsstrahlung emission.
@@ C.2.1 Comparison in the deceleration region
As shown in the figures, the profiles of $\Gamma$ exhibit very good agreement between the analytical model and the simulation results. When excluding the far upstream region ($\Gamma\approx\Gamma_{u}$) where deceleration is negligible, broad agreement is also observed for the profiles of $\hat{T}$ and $x_{l}$. The discrepancies in $\hat{T}$ and $x_{l}$ between the analytical model and the simulation in the deceleration region remain within a factor of 2 for $\Gamma_{u}=6$ and $10$, and at most $\sim 3$ for $\Gamma_{u}=2$. Given the simplified assumptions of the analytical model, this level of agreement highlights its ability to capture the essential features of shock dissipation physics. Notably, it is remarkable that even for the mildly relativistic shock with $\Gamma_{u}=2$, the model exhibits reasonable agreement. This suggests that, although the model is derived under the assumption that the flow is highly relativistic ($\Gamma\gg 1$), it can be extended to the mildly relativistic regime, at least down to $\Gamma_{u}\sim 2$.
@@
Compared to the mildly relativitic shock $\Gamma_{u}=2$, $\Gamma_{u}=6$ and $\Gamma_{u}=10$ shocks exhibits better agreement in the deceleration region as mentioned earlier. The temperature profile shows a particularly good match with the analytical model, as seen in the figures. The quantity $x_{l}$ also shows good agreement between the simulation and the analytical solution. Regarding the spatial distribution of $x_{l}$, a particularly good match is observed as the flow approaches the downstream region. However, a gap between the analytical model and the simulation gradually develops toward the upstream direction. This discrepancy arises because the advected photons in the simulation include not only scattered-back photons but also those generated by thermal bremsstrahlung. The thermal bremsstrahlung process produces numerous low-energy photons ($\lesssim m_{e}c^{2}$), which do not contribute to $\gamma$-$\gamma$ pair production. As a result, $x_{l}$ in the simulation accounts for a population of photons that do not contribute to the opacity.
It is worth noting that the qualitative difference between the mildly relativistic shock ($\Gamma_{u}=2$) and highly relativistic shocks ($\Gamma_{u}=6$ and $10$) is clearly reflected in a systematic difference in the values of the free parameters. Specifically, while the free parameters for highly relativistic shocks remain consistently within a certain range, the parameters for the mildly relativistic shock systematically exhibit lower values. Consequently, the shock width $\Delta\tilde{\tau}$ follows the expected scaling law, $\Delta\tilde{\tau}\propto\Gamma_{u}^{3}$, for $\Gamma_{u}\geq 6$, as reported in ILN20a, whereas for $\Gamma_{u}=2$, a noticeable deviation from this scaling has been observed, as also pointed out in ILN20a.

### v2

@@ 8 Summary and conlusions
In this study, we performed Monte Carlo simulations of photon starved relativistic RRMS, incorporating the effect of photon escape from the upstream region under the assumption of steady-state flow in a planar geometry. These simulations are applicable to RRMS propagating through shallowly decaying density profiles, such as those found in stellar winds. The calculations were carried out for shock Lorentz factors of $\Gamma_{u}=2$, $3.5$, 66, $10$, and $15$ marking the first simulation to model relativitic shock breakout emission from the first principles in the photon starved regime. [^13] Lundman and Beloborodov (2021) conducted a first principles calculation of relativistic shock breakout in the “photon rich” regime, where photon production and pair production can be neglected. We explored a wide range of energy fractions carried by escaping photons, $f_{esc}$, and analyzed the resulting shock structures and the spectral properties of the escaping radiation. We also conducted a detailed comparison between our numerical results and the shock structures predicted by an analytical model (GNL18), identifying deviations that suggest the need to reconsider the predicted observables of relativistic shock breakouts in a wind environment.
@@
    While our results confirm the overall trends predicted by the analytical model, they also reveal systematic deviations. The model successfully reproduces the structure of infinite shocks for all tested Lorentz factors ($\Gamma_{u}=2$, $3.5$, 66, and $10$), and captures finite shock structures well for $\Gamma_{u}=2$ across a wide range of $f_{\rm esc}$. However, for higher Lorentz factors ($\Gamma_{u}=3.5$, $6$, $10$, and $15$), our simulations yield systematically narrower shock widths than predicted, particularly at large $f_{\rm esc}$. Particularly, we find a scaling of $\Delta\tilde{\tau}\propto f_{\rm esc}^{-1.9}$, compared to the analytical prediction of $\Delta\tilde{\tau}\propto f_{\rm esc}^{-1}$. Furthermore, while the analytical model predicts a positive scaling with Lorentz factor ($\Delta\tilde{\tau}\propto\Gamma_{u}$), our simulations show an inverse trend, $\Delta\tilde{\tau}\propto\Gamma_{u}^{-1.8}$. These discrepancies arise from a shift in the dominant deceleration mechanism at high escape fractions. In this regime, $\gamma$$\gamma$ pair production between high-energy photons—upscattered by subshock-heated pairs—and the bulk population of counterstreaming photons becomes the main source of flow deceleration, overtaking the interaction between counterstreaming photons and the advected plasma originally assumed in the analytical model.
@@
    Due to efficient temperature regulation in the downstream region, the spectra of escaping photons consistently peak at $E_{p}\approx 300$-$600~{\rm keV}$ in the shock frame, largely independent of $f_{\rm esc}$ and $\Gamma_{u}$. The spectral shape deviates significantly from a Wien distribution, exhibiting broadening at both low and high energies. At lower energies ($h\nu\ll E_{p}$), a flat spectral component with $f_{\nu}\propto\nu^{0}$ emerges due to free-free emission, and its prominence increases with larger $f_{esc}$. At higher energies ($h\nu>E_{p}$), our simulations reveal a spectral extension, particularly for $\Gamma_{u}=6$, $10$, and $15$, originating from inverse Compton scattering by hot pairs in the post-subshock region. The strength and extent of this high-energy component increase with both $\Gamma_{u}$—reflecting higher post-subshock temperatures—and $f_{esc}$—due to reduced optical depth for $\gamma$$\gamma$ pair production. These results indicate that the spectrum undergoes rapid broadening during breakout as $f_{\rm esc}$ increases.
@@
    Based on the deviations from the analytical model, we derive an updated closure relation applicable to relativistic breakouts with $\Gamma_{bo}\gtrsim 6$. We also present revised estimates of breakout observables in wind environments using a spherical explosion model. Notably, we find that the breakout duration becomes largely insensitive to the explosion energy and ejecta mass, depending primarily on the progenitor radius and wind optical depth, following $t_{bo}\sim 3.8\times 10^{2}E_{53}^{-0.15}M_{\rm ej,5}^{0.1}R_{*,11}^{1.1}\tau_{w*}^{1.1}$. This duration is significantly longer—by orders of magnitude—than that predicted by the analytical model for the same explosion parameters, reflecting the substantially reduced shock width in finite shocks. For example, a highly energetic explosion with isotropic equivalent energy $E_{ex}\sim 5\times 10^{53}{\rm erg}$ from a compact progenitor ($R_{*}\sim 10^{11}{\rm cm}$) surrounded by a wind of optical depth $\tau_{*}\sim 1$ can produce a relativistic breakout with $\Gamma_{bo}\sim 6$. The resulting emission peaks at $\sim 3kT_{obs}\sim 3~{\rm MeV}$, with total radiated energy $E_{bo}\sim 10^{50}~{\rm erg}$ and a duration of approximately $300~{\rm s}$. In contrast, the analytical model predicts a sub-second duration for the same parameters. Such MeV-range emission may be detectable by current gamma-ray observatories such as Fermi-GBM up to distances of $\sim$100 Mpc. Future missions with enhanced sensitivity in the MeV band, including AMEGO-X, COSI, and e-ASTROGAM, could extend this detection horizon to several hundred Mpc, significantly increasing the prospects for observing relativistic shock breakouts.
With the limitations summarized in Section 7 in mind, the simulations presented here are, to the best of our knowledge, among the most sophisticated to date for RRMS breakout in wind-like environments. Our calculations, however, cover only a limited region of the parameter space in $\Gamma_{u}$ and $f_{\rm esc}$, leaving some uncertainty in our rough estimates of the breakout emission. Ultimately, fully dynamical simulations will be required to validate and extend these results. In this context, the present work should be regarded as a first step and a useful baseline for interpreting observations and for guiding future time-dependent studies.
@@ Acknowledgements
This work was supported by JSPS KAKENHI Grant Number JP24H01826, JP23K25874, and JP25H00675. S.N. is supported by the JST ASPIRE Program ”RIKEN-Berkeley mathematical quantum science initiative”. AL and EN acknowledge support by the Israel Science Foundation grant 1995/21, and a BSF grant (2022314) and ERC grant JetNS (818899). Numerical computations and data analyses were carried out on the supercomputer HOKUSAI BigWaterfall2 (HBW2) at RIKEN, Cray XC50 and HPE Cray XD2000 at the Center for Computational Astrophysics, National Astronomical Observatory of Japan, and Yukawa-21 at the Yukawa Institute for Theoretical Physics, Kyoto University.
Data Availability
@@ Appendix A Spectrum at low energies
Figure 8: Shock-frame, $f_{\nu}$ flux of escaping photons normalized by the total kinetic energy flux of baryons at the upstream boundary, $F_{b}=\Gamma_{u}(\Gamma_{u}-1)n_{u}m_{p}c^{3}\beta_{u}$. Each panel display the results for shock velocities $\Gamma_{u}=2$ (top left), $3.5$ (top right), 66 (bottom left), and $10$ (bottom right), respectively. Within each panel, various lines represent different escape fractions, as specified in the legend. Note that the energy range of the spectra varies across different panels.
@@
Figure 9: The low cutoff energy for free-free emission, estimated from Equation (15), is shown for $\Gamma_{u}=2$ (top left), $\Gamma_{u}=3.5$ (top right), $\Gamma_{u}=6$ (bottom left), and $\Gamma_{u}=10$ (bottom right). Within each panel, various lines represent different escape fractions, as specified in the legend.
As the equation indicates, $\nu_{cut}$ varies with position, depending on the local density of pairs and temperature. In the post-subshock region, where most photons are produced, the cutoff frequency progressively decreases in the downstream direction. The spatial profile of $\nu_{cut}$ is displayed in Figure 9. As seen in the figure, $h\nu_{cut}$ reaches maximum values of approximately $0.4$, 22, 88, and $30~{\rm eV}$ in simulations with the highest escape fractions for $\Gamma_{u}=2$, $3.5$, 66, and $10$, respectively, at the immediate post-subshock region. These peak values are higher for higher escape fractions, reflecting elevated post-subshock temperatures. Beyond the peak, $h\nu_{cut}$ decreases to about $0.03$, $0.7$, $0.2$, and $0.3~{\rm eV}$ at $\tau_{*}=10$, which corresponds to roughly one diffusion length downstream from the subshock. While the maximum values somewhat depend on the escape fractions particularly for high $\Gamma_{u}$ shocks, the profile beyond $\tau_{*}\gtrsim 1$ shows little variation across different escape fractions. This stability is attributed to the minimal variation in the temperature profile that follows the rapid cooling of subshock-heated pairs (Figure 1-4).
@@ C.1 Conversion of Optical Depth
As shown in the previous section, the analytical model provides the shock profile as a function of the net optical depth, $\tau$, without explicitly specifying its relation to the spatial scale. Consequently, converting $\tau$ into a physical length scale is necessary to facilitate the comparison between the simulation and the model. In this study, we achieve this by translating $\tau$ to the pair unloaded Thomson optical depth $\tilde{\tau}$ ($2\Gamma_{u}\sigma_{T}z$), following the methodology outlined in GNL18:
@@
[…]
represents an approximate value for the typical photon energy of the counterstreaming photons observed in the rest frame of the pairs, normalized by $m_{e}c^{2}$ in the case of Compton scattering. At the same time, it also represents the product of the counterstreaming photon energy and the thermal photon energy advected with the flow, both evaluated in the comoving frame. [^17] When $X\gg 1$ and the typical advected photon energy satisfies $kT\gg m_{e}c^{2}$, Equation (24) provides a good approximation for the $\gamma$-$\gamma$ pair production cross-section. There is a slight modification to the cross-section given in Equation (24) compared to GNL18, which does not include the additional +0.5+0.5 term in the numerator. This term is added solely to improve accuracy in the deep KN regime ($X\gg 1$), although it is not crucial, as no significant differences are observed regardless of which prescription is used. The parameter $a$, which is expected to be of order unity, is treated as an additional free parameter of the analytical model, along with $\eta$, both of which depend on the energy and angular distributions of pairs and photons. The above prescription approximates that each particle in the primary beam contributes equally to the opacity for the counterstreaming photons, regardless of whether the particle is a pair or a photon.
@@ C.2 Fitting to the simulation results
[…]
@@ C.2.1 Comparison in the deceleration region
As shown in the figures, the profiles of $\Gamma$ exhibit very good agreement between the analytical model and the simulation results. When excluding the far upstream region ($\Gamma\approx\Gamma_{u}$) where deceleration is negligible, broad agreement is also observed for the profiles of $\hat{T}$ and $x_{l}$. The discrepancies in $\hat{T}$ and $x_{l}$ between the analytical model and the simulation in the deceleration region remain within a factor of $\sim 2$ for $\Gamma_{u}=3.5$, $6$, and $10$, and at most $\sim 3$ for $\Gamma_{u}=2$. Given the simplified assumptions of the analytical model, this level of agreement highlights its ability to capture the essential features of shock dissipation physics. Notably, it is remarkable that even for the mildly relativistic shock with $\Gamma_{u}=2$, the model exhibits reasonable agreement. This suggests that, although the model is derived under the assumption that the flow is highly relativistic ($\Gamma\gg 1$), it can be extended to the mildly relativistic regime, at least down to $\Gamma_{u}\sim 2$.
@@
Compared with the $\Gamma_{u}=2$ case, shocks with higher Lorentz factors ($\Gamma_{u}=3.5$, $6$, and $10$) exhibit better agreement in the deceleration region, as mentioned earlier. The temperature profile shows a particularly good match with the analytical model, as seen in the figures. The quantity $x_{l}$ also shows good agreement between the simulations and the analytical solution, particularly for the high Lorentz factor shocks ($\Gamma_{u}=6$ and $10$). For $\Gamma_{u}=3.5$, as in the $\Gamma_{u}=2$ case, the simulated pair fraction $n_{\pm}/n$ is in closer agreement with the analytical $x_{l}$ than is the simulated $x_{l}$. Nonetheless, the simulated $x_{l}$ at $\Gamma_{u}=3.5$ exhibits slightly better agreement with the analytical $x_{l}$ than at $\Gamma_{u}=2$, indicating behavior intermediate between the low- and high-$\Gamma_{u}$ regimes. For $\Gamma_{u}=6$ and $10$, the spatial profile of $x_{l}$ shows particularly good agreement between the simulations and the analytical solution, especially toward the downstream region. However, a gap between the analytical model and the simulation gradually develops toward the upstream direction. This discrepancy arises because the advected photons in the simulation include not only scattered-back photons but also those generated by thermal bremsstrahlung. The thermal bremsstrahlung process produces numerous low-energy photons ($\lesssim m_{e}c^{2}$), which do not contribute to $\gamma$-$\gamma$ pair production. As a result, $x_{l}$ in the simulation accounts for a population of photons that do not contribute to the opacity.
It is worth noting that the qualitative difference between the mildly relativistic shocks ($\Gamma_{u}=2$ and $3.5$) and highly relativistic shocks ($\Gamma_{u}=6$ and $10$) is clearly reflected in a systematic difference in the values of the free parameters. Specifically, while the free parameters for highly relativistic shocks remain consistently within a certain range, the parameters for the mildly relativistic shock systematically exhibit lower values. Consequently, the shock width $\Delta\tilde{\tau}$ follows the expected scaling law, $\Delta\tilde{\tau}\propto\Gamma_{u}^{3}$, for $\Gamma_{u}\geq 6$, as reported in ILN20a, whereas at lower Lorentz factor $\Gamma_{u}\lesssim 3.5$, a noticeable deviation from this scaling emerges, as also pointed out in ILN20a.